Find the equation to the straight line cutting of an intercept unity fro the positive directio of Y axis and inclined at 45to the X axis

Grade:10

1 Answers

Vikas TU
4 years ago
The equation is:
y-y1= m(x-x1)
Since the angle for the line is 45, then :
m= tan45 = 1
Since it cuts the y axis at +1 , then the point (0,1) belongs to the line:
==> y-1 = 1*(x-0)
==> y= x + 1